# Re-visiting Skip-Gram Negative Sampling: Dimension Regularization for More Efficient Dissimilarity Preservation in Graph Embeddings

### v1

@@ Abstract.
A wide range of graph embedding objectives decompose into two components: one that attracts the embeddings of nodes that are perceived as similar, and another that repels embeddings of nodes that are perceived as dissimilar. Because real-world graphs are sparse and the number of dissimilar pairs grows quadratically with the number of nodes, Skip-Gram Negative Sampling (SGNS) has emerged as a popular and efficient repulsion approach. SGNS repels each node from a sample of dissimilar nodes, as opposed to all dissimilar nodes. In this work, we show that node-wise repulsion is, in aggregate, an approximate re-centering of the node embedding dimensions. Such dimension operations are much more scalable than node operations. The dimension approach, in addition to being more efficient, yields a simpler geometric interpretation of the repulsion. Our result extends findings from the self-supervised learning literature to the skip-gram model, establishing a connection between skip-gram node contrast and dimension regularization. We show that in the limit of large graphs, under mild regularity conditions, the original node repulsion objective converges to optimization with dimension regularization. We use this observation to propose an algorithm augmentation framework that speeds up any existing algorithm, supervised or unsupervised, using SGNS. The framework prioritizes node attraction and replaces SGNS with dimension regularization. We instantiate this generic framework for LINE and node2vec and show that the augmented algorithms preserve downstream performance while dramatically increasing efficiency.
@@ 1. Introduction
Graph embedding algorithms use the structure of graphs to learn node-level embeddings. Across unsupervised and supervised graph embedding algorithms, their loss functions serve the two roles of preserving similarity and dissimilarity. Nodes that are similar in the input graph should have similar embeddings, while dissimilar nodes should have dissimilar embeddings (Böhm et al., 2022). The push and pull of the similarity and dissimilarity objectives are key: in the absence of a dissimilarity objective, the loss would be minimized by embedding all nodes at a single embedding point, a degenerate and useless embedding. Often, enforcing dissimilarity is much more expensive than similarity, owing to the generally sparse nature of graphs and the number of pairs of dissimilar nodes growing quadratically with the size of the graph. Enforcing dissimilarity is also complex for graphs because graph data frequently have missing edges or noise (Young et al., 2021; Newman, 2018). In this paper, we show that while many past works have focused on explicitly repelling pairs of dissimilar nodes, the repulsion can be replaced with a regularization of the embedding dimensions. The dimension-based approach improves both scalability and representation quality.
@@
The skip-gram (SG) model is one of the most popular approaches to graph embeddings (Ahmed et al., 2013; Yang et al., 2024) and can be decomposed into preserving similarity and dissimilarity. Further, skip-gram negative sampling (SGNS) is the dominant method to efficiently approximate dissimilarity preservation. Instead of repelling all pairs of dissimilar nodes, SGNS repels only a sample of dissimilar nodes per pair of similar nodes. SGNS is utilized in LINE (Tang et al., 2015) and node2vec (Grover and Leskovec, 2016), for instance, and has been shown to yield strong downstream performance in practice. However, several analytical issues with SGNS have also been identified. First, SGNS approximates the original loss function but also introduces a bias by re-scaling the relative importance of preserving similarity and dissimilarity (Rudolph et al., 2016). Second, with SGNS, in the limit as the number of nodes in the graph approaches infinity, the similarities among embeddings diverge from the similarities among nodes in the graph (Davison and Austern, 2023). Although SGNS has been used to learn both graph and word embeddings (Mikolov et al., 2013; Mimno and Thompson, 2017), we focus on the graph context because, for graph embeddings in particular, SGNS remains a popular method for preserving dissimilarity (Chami et al., 2022).
@@
In this paper, we propose a change in perspective and show that node repulsion in the SG model can be achieved via dimension centering. We draw inspiration from recent advances in the self-supervised learning (SSL) literature, which show an equivalence between sample-contrastive learning and dimension-contrastive learning (Garrido et al., 2022; Bardes et al., 2022). Sample-contrastive learning explicitly repels dissimilar pairs while dimension-contrastive learning repels the dimensions from each other. The findings from the SSL literature allow us to directly re-interpret the spectral graph embedding loss function, for example, as achieving dissimilarity by imposing orthogonality among the embedding dimensions. In the case of spectral embeddings, repelling all pairs of nodes is exactly equal to repelling all pairs of dimensions, which is much more scalable given that the number of dimensions is much less than the number of nodes.
@@
The known parallels between sample and dimension contrast, however, do not suggest whether SG loss functions can be also re-interpreted from the dimension perspective. In this paper, we begin by characterizing the degenerate embedding behavior when the dissimilarity term is removed altogether. We prove that, under mild initialization conditions, when only positive edges are considered, the embeddings collapse into a lower dimensional space, which also commonly occurs in self-supervised learning (Jing et al., 2022). However, as the dimensions approach collapse, the dissimilarity term also approaches a dimension regularizer that centers the embeddings around the origin. Our findings show that while the dissimilarity term in the SG loss is not itself a dimension regularizer, when the term is most needed to counteract the similarity attraction, dissimilarity preservation can be achieved via regularization.
@@
We operationalize the dimension-based replacement with an algorithm augmentation. We augment existing algorithms using SGNS by making two modifications. First, the augmentation prioritizes the similarity-preservation component, disregarding dissimilarity when possible. This is also desirable because, in real-world graph data, the lack of similarity between two nodes does not necessarily suggest the two nodes are dissimilar; it is also possible that data are missing or noisy (Newman, 2018; Young et al., 2021). Second, when the embeddings begin to collapse after optimizing only for similarity preservation for a fixed number of epochs, our augmentation repels nodes from each other using a dimension regularizer that aims to re-center the embeddings about the origin.
@@
    In Section 2, we introduce a framework that maps node repulsion to dimension regularization. We show that instead of shortcutting the full skip-gram loss function with SGNS and repelling a sample of pairs, the negative function can be approximated with a dimension regularization inducing origin-centered embeddings. We prove that as the need for node repulsion grows, optimizing the regularizer converges to optimizing the skip-gram loss. This framework extends the equivalence between sample-contrastive objectives and dimension regularization established in self-supervised learning.
@@
    In Section 3 we introduce a generic algorithm augmentation that replaces SGNS with dimension regularization for any existing skip-gram embedding algorithm. We instantiate the augmentation for node2vec and LINE, reducing the repulsion complexity from $\mathcal{O}(n)$ to $\mathcal{O}(d)$ per epoch.
@@
    In Section 4 we present empirical evaluations of our augmented LINE and node2vec algorithms. Our results show that replacing SGNS with dimension regularization greatly reduces runtime while also improving performance on transductive link prediction. We also show that dimension regularization especially outperforms baselines when graph connectivity is high in a synthetic example, when the blocks in a stochastic block model (SBM) are less distinguishable.
@@ 2. From Node Repulsion to Dimension Regularization
In this section, we introduce our loss decomposition framework where a function $P$ operationalizes similarity preservation and a “negative” function $N$ achieves dissimilarity preservation. We then show that instead of optimizing negative functions with costly node repulsions, we can instead optimize via dimension regularization. Crucially, we show in Subsection 2.2 we show that when node repulsion is needed, the negative function in the skip-gram loss can be optimized via dimension regularization.
@@
Symbol Meaning
$G,V,E$ Graph $G$ with vertices $V$ and edges $E$
$n,m$ number of nodes and edges respectively
$d$ number of embedding dimensions
$N$, $P$ negative and positive loss functions
$S$ similarity matrix $\in\mathbb{R}^{n\times n}$
$X$ node embedding matrix $\in\mathbb{R}^{n\times d}$
$X_{i}$ $i^{th}$ row of $X$, as a column vector
$X_{.j}$ $j^{th}$ column of $X$, as a column vector
$P_{\alpha}$ Probability distribution with parameter $\alpha$
$k$ number of negative samples per positive update
$\eta$ learning rate
$D_{x}$ diagonal matrix where $x$ is the diagonal
$\mathcal{C}$ the constriction of the embeddings (Def. 2.2)
$\vec{1},\mathbf{1}$ a vector and matrix of all ones, respectively.
@@
The decomposition in Equation (1) applies to nearly all unsupervised graph embedding objectives as well as many supervised learning objectives, where supervision is provided in the form of node labels. In the recent graph embedding survey by Chami et al. (2022), the decomposition applies to all unsupervised methods except for Graph Factorization (Ahmed et al., 2013), which does not include a negative function $N$. Examples of popular decomposable loss functions are matrix reconstruction error (e.g., spectral embeddings) as well as softmax (e.g. node2vec (Grover and Leskovec, 2016) and LINE (Tang et al., 2015)). The decomposition also applies to supervised methods that regularize for graph structure ($\beta>0$ as defined in Chami et al. (2022)), such as Neural Graph Machines (Bui et al., 2018) and Planetoid (Yang et al., 2016).
@@
In the context of graphs, performing gradient descent on the negative example loss $N$ is a costly operation as the gradient repels all pairs of nodes, resulting in $\mathcal{O}\left(n^{2}\right)$ time complexity. In this paper, we build on the argument that the costly node-wise operation can be replaced with a more efficient dimension-wise operation. Optimizing from the dimension perspective also yields a geometric interpretation. This interpretation is illustrated in Figure 1.
@@
Figure 1. Nearly all unsupervised and many supervised graph embedding loss functions define a “negative” function that repels embeddings of dissimilar nodes. We show that instead of repelling pairs of nodes (orange), which is costly, the negative function in the popular skip-gram (SG) loss can be approximated with a dimension regularizer. The regularizer penalizes non-zero dimension means and is efficient given that $d\ll n$. This result complements the duality between squared-loss and dimension covariance established in the self-supervised learning literature.
@@
Below, we will map the two dominant negative functions found in graph embedding algorithms to dimension regularizations. The two loss functions are spectral loss functions and skip-gram loss functions. The mapping of spectral loss functions to dimension covariance regularization directly ports over a result from the self-supervised learning literature on non-contrastive learning (Bardes et al., 2022; Garrido et al., 2022). Our novel contribution is a mapping from the skip-gram loss to a regularizer that induces origin-centered dimensions. Proof for all propositions below are included in Appendix A.

### v2

# Bypassing Skip-Gram Negative Sampling: Dimension Regularization as a More Efficient Alternative for Graph Embeddings

@@ Abstract.
A wide range of graph embedding objectives decompose into two components: one that enforces similarity, attracting the embeddings of nodes that are perceived as similar, and another that enforces dissimilarity, repelling the embeddings of nodes that are perceived as dissimilar. Without repulsion, the embeddings would collapse into trivial solutions. Skip-Gram Negative Sampling (SGNS) is a popular and efficient repulsion approach that prevents collapse by repelling each node from a sample of dissimilar nodes. In this work, we show that when repulsion is most needed and the embeddings approach collapse, SGNS node-wise repulsion is, in the aggregate, an approximate re-centering of the node embedding dimensions. Such dimension operations are more scalable than node operations and produce a simpler geometric interpretation of the repulsion. Our theoretical result establishes dimension regularization as an effective and more efficient, compared to skip-gram node contrast, approach to enforcing dissimilarity among embeddings of nodes. We use this result to propose a flexible algorithm augmentation framework that improves the scalability of any existing algorithm using SGNS. The framework prioritizes node attraction and replaces SGNS with dimension regularization. We instantiate this generic framework for LINE and node2vec and show that the augmented algorithms preserve downstream link-prediction performance while reducing GPU memory usage by up to $33.3\%$ and training time by $23.4\%$. Moreover, we show that completely removing repulsion (a special case of our augmentation framework) in LINE reduces training time by $70.9\%$ on average, while increasing link prediction performance, especially for graphs that are globally sparse but locally dense. Global sparsity slows down dimensional collapse, while local density ensures that node attraction brings the nodes near their neighbors. In general, however, repulsion is needed, and dimension regularization provides an efficient alternative to SGNS.
journalyear: 2025copyright: cc
@@ 1. Introduction
Graph embedding algorithms use the structure of graphs to learn node-level embeddings. In unsupervised and supervised graph embedding algorithms, loss functions attempt to preserve similarity and dissimilarity. Nodes that are similar in the input graph should have similar embeddings, while dissimilar nodes should have dissimilar embeddings (Böhm et al., 2022; Tsitsulin et al., 2018). The push and pull of the similarity and dissimilarity objectives are key: in the absence of a dissimilarity objective, the loss would be minimized by embedding all nodes at a single embedding point, a degenerate and useless embedding. Often, enforcing dissimilarity is much more expensive than similarity, owing to the generally sparse nature of graphs and the number of pairs of dissimilar nodes growing quadratically with the size of the graph. Enforcing dissimilarity is also complex for graphs because graph data frequently have missing edges or noise (Young et al., 2021; Newman, 2018). Here, we show that while many past works have focused on repelling pairs of dissimilar nodes, the repulsion can be replaced with a regularization of the embedding dimensions, which is more scalable.
@@
The skip-gram (SG) model is one of the most popular approaches to graph embeddings (Ahmed et al., 2013; Yang et al., 2024). In the SG model, a target node is similar to a source node if it is in the source node’s neighborhood. Furthermore, skip-gram negative sampling (SGNS) is a dominant method to efficiently approximate dissimilarity preservation. Instead of repelling all pairs of dissimilar nodes, SGNS repels only a sample of dissimilar nodes per pair of similar nodes. SGNS is utilized in LINE (Tang et al., 2015) and node2vec (Grover and Leskovec, 2016), for instance, and has been shown to yield strong downstream performance. However, several analytical issues with SGNS have also been identified. First, SGNS introduces a bias by re-scaling the relative importance of preserving similarity and dissimilarity (Rudolph et al., 2016). Second, with SGNS, in the limit as the number of nodes in the graph approaches infinity, the similarities among embeddings diverge from the similarities among nodes in the graph (Davison and Austern, 2023). Although SGNS has been used to learn both graph and word embeddings (Mikolov et al., 2013; Mimno and Thompson, 2017), we focus on the graph context because, for graph embeddings in particular, SGNS remains a popular method for preserving dissimilarity (Chami et al., 2022).
@@
In this paper, we propose a change in perspective and show that node repulsion in the SG model can be achieved via dimension centering. If $X$ is an embedding matrix where the rows are node embeddings, “dimensions” refers to the columns of $X$. We draw inspiration from advances in the self-supervised learning (SSL) literature, which show an equivalence between sample-contrastive learning and dimension-contrastive learning (Garrido et al., 2022; Bardes et al., 2022). Sample-contrastive learning explicitly repels dissimilar pairs while dimension-contrastive learning regularizes dimensions.
@@
The known parallels between sample and dimension contrast, however, do not suggest whether SG loss functions can also be re-interpreted from the dimension perspective. We present a novel theoretical contribution that extends the dimension-contrastive approach to a new class of loss functions. Our findings show that while the dissimilarity term in the SG loss is not itself a dimension regularizer when the term is most needed to counteract the attraction of similarity, the dissimilarity preservation can be achieved via regularization. We begin by characterizing the degenerate embedding behavior when the dissimilarity term is removed altogether. We prove that, under initialization conditions, [^1] Our asymptotic result holds when the learning rate approaches zero and the embedding initializations approach the origin, an assumption inherited from the implicit regularization in matrix factorization established in Gunasekar et al. (2017). when only positive pairs are considered, the embeddings collapse into a low-dimensional space. However, as the dimensions approach collapse, the dissimilarity term approaches a dimension-mean regularizer.
@@
We operationalize the dimension-based approach with an algorithm augmentation framework. We augment existing algorithms using SGNS by making two modifications. First, the augmentation framework prioritizes similarity preservation over dissimilarity preservation. This is desirable because, in real-world graph data, the lack of similarity between two nodes does not necessarily suggest that the two nodes are dissimilar. Second, when the embeddings begin to collapse after optimizing only for similarity preservation for a fixed number of epochs, our augmentation framework repels nodes from each other using a regularizer that induces embedding dimensions centered on the origin.
@@
    In Section 3, we map node repulsion to dimension regularization for the SG model. We show that instead of shortcutting the full skip-gram loss function with SGNS and repelling a sample of pairs, the repulsion function can be approximated with a dimension-mean regularization. We prove that as the need for node repulsion grows, optimizing the regularizer converges to optimizing the skip-gram loss. Our result extends the equivalence between sample-contrastive objectives and dimension regularization established in self-supervised learning to a new class of loss functions.
@@
    In Section 4, we introduce a generic algorithm augmentation framework that prioritizes node attraction and replaces SGNS with occasional dimension regularization for any existing SG algorithm. We instantiate the augmentation framework for node2vec and LINE, reducing the repulsion complexity from $\mathcal{O}(n)$ to $\mathcal{O}(d)$ vector additions per epoch. (The number of embedding dimensions $d$ is much less than the number of nodes $n$.)
@@
    In Section 5, we show that our augmentation framework reduces runtime and memory usage while also preserving link-prediction performance. Moreover, in globally sparse networks with high local density, removing repulsion altogether, a special case of our framework, even improves performance for LINE. However, in networks with low local density, repulsion is needed, and dimension regularization provides an efficient solution.
2. Related Works
In this section, we review the popular use of SGNS within graph embeddings as well as its limitations. Our approach is also similar in spirit to the growing body of literature on non-contrastive learning and can be used in concert with existing scalable graph learning methods.
2.1. Skip-Gram Negative Sampling
SGNS was introduced in word2vec by Mikolov et al. (2013) as an efficient method for learning word embeddings. While the softmax normalization constant is costly to optimize, Mikolov et al. (2013) modeled SGNS after Noise Contrastive Estimation (NCE) (Gutmann and Hyvärinen, 2012), which learns to separate positive samples from samples drawn from a noise distribution. SGNS has since been adopted for graph representation learning, where it is utilized in both unsupervised (Grover and Leskovec, 2016; Tang et al., 2015; Perozzi et al., 2014) and supervised (Yang et al., 2016) skip-gram models.
At the same time, there are many known limitations of SGNS. Rudolph et al. (2016) place SGNS embeddings within the framework of Exponential Family Embeddings and note that SGNS downweights the magnitude of the negative update and leads to biased embeddings, relative to the gradients of the non-sampled objective. Second, Davison and Austern (2023) examine the limiting distribution of embeddings learned via SGNS and show that the distribution decouples from the true sampling distribution in the limit. Last, it has also been shown that the optimal noise distribution and the corresponding parameters can vary by dataset (Yang et al., 2020).
2.2. Non-contrastive Self-Supervised Learning
Energy-Based Models in self-supervised learning are a unified framework for balancing similarity and dissimilarity (LeCun et al., 2007). As in our decomposition, energy-based models ensure that similar pairs have low energy and dissimilar pairs have high energy. Within energy-based models, there has been more focus across both vision and graph representation learning on contrastive models, which explicitly repel dissimilar pairs (In et al., 2023; Li et al., 2023; Yang et al., 2023; Zhang et al., 2023). However, given the computational complexities of pairwise contrast, there is a growing body of work on non-contrastive representation learning methods, which do not use negative samples. Embedding collapse is the main challenge facing non-contrastive methods, and several mitigation strategies have emerged, such as asymmetric encoders in SimSiam (Chen and He, 2021), momentum encoders in BYOL (Grill et al., 2020), and redundancy reduction in Barlow Twins (Zbontar et al., 2021). Garrido et al. (2022) establish a duality between dimension-regularization based non-contrastive approaches and standard contrastive learning; however, the work specifically analyzes the squared-loss term, and there are no existing works, to our knowledge, establishing a connection between skip-gram loss and non-contrastive methods.
2.3. Scalable and Efficient Graph Learning
Past works on developing scalable and efficient representation learning for large graphs have focused on data-driven scalability, utilizing node/edge sampling (Hamilton et al., 2017; Chen et al., 2018; Zeng et al., 2020), graph partitioning (Chiang et al., 2019; Zhu et al., 2019), and graph coarsening (Liang et al., 2021). The sampling-based approaches mitigate the exponential growth of neighbors as the depth of message-passing layers increases in GNNs. One approach is to sample at the layer level, such as in GraphSAGE (Hamilton et al., 2017) and FastGCN (Chen et al., 2018), where the neighbors of each node are sampled in each layer. An alternative approach, such as GraphSAINT (Zeng et al., 2020), is to first sample a subgraph and then train a full GNN, without layer sampling, on the subgraph. Likewise, partitioning-based approaches also prevent neighborhood explosion. For instance, ClusterGCN (Chiang et al., 2019) performs message passing within clusters to reduce neighborhood scope, while GraphVite (Zhu et al., 2019) leverages partitioning to train in parallel. Finally, coarsening approaches leverage the multi-scale nature of graphs. For example, MILE (Liang et al., 2021) first learns embeddings for the structural backbone of a graph, which is much smaller, before refining the backbone embeddings onto the full, original graph. Our approach is orthogonal to data-driven approaches to scalability and can be used in conjunction; for example, our dimensionality-regularization bypass improves the efficiency of learning skip-gram embeddings for graph backbones or cluster partitions.
Instead, our approach is more similar to scalability approaches that propose efficiency bypasses for loss functions or optimization processes. For instance, Zhu et al. (2025) identify that the outputs of normalization layers in transformers approximate the hyperbolic tangent function, and replacing normalization layers with element-wise $\tanh$ transformations preserves accuracy while reducing training time. Meanwhile, Loveland et al. (2025) observe that one purpose of negative sampling in collaborative filtering is to increase the rank of the embedding matrix; the authors show that replacing negative sampling with a regularization on matrix rank, which is more efficient, effectively reduces compute cost. In contrast to these two prior works proposing bypasses, our work contributes a theoretical result showing that as the number of nodes increases, the difference between the expensive operation, SGNS, and the efficiency bypass, dimension-mean regularization, vanishes.
@@ 3. From Node Repulsion to Dimension Regularization
In this section, we introduce a generic loss decomposition where a function $P$ operationalizes similarity preservation and a “negative” function $N$ achieves dissimilarity preservation. We then show that instead of optimizing negative functions with costly node repulsions, we can instead regularize dimensions. Crucially, in Section 3.2, we show that when node repulsion is needed, the negative function in the skip-gram loss can be optimized via dimension regularization.
@@
[…]
[Notation table]One column of symbols on the left and another column with the corresponding meanings on the right. Symbol Meaning $G,V,E$ Graph $G$ with nodes $V$ and edges $E$ $n,m$ number of nodes and edges respectively $d$ number of embedding dimensions $N$, $P$ negative and positive loss functions $S$ similarity matrix $\in\{0,1\}^{n\times n}$ $X$ node embedding matrix $\in\mathbb{R}^{n\times d}$ $X_{i}$ $i^{th}$ row of $X$, as a column vector $X_{.j}$ $j^{th}$ column of $X$, as a column vector $P_{\alpha}$ Probability distribution with parameter $\alpha$ $p,q$ node2vec random-walk bias parameters (Grover and Leskovec, 2016) $k$ number of negative samples per positive pair $b$ average number of positive pairs per node $\eta$ learning rate $D_{x}$ diagonal matrix where $x$ is the diagonal $\mathcal{C}$ the constriction of the embeddings (Def. 3.2) $\vec{1},\mathbf{1}$ a vector and a matrix of all ones, respectively $[d]$$\{1,2,\dots,d\}$
@@
The decomposition in Eq. (1) applies to nearly all unsupervised graph embedding objectives as well as many supervised learning objectives, where supervision is provided in the form of node labels. In the graph embedding survey by Chami et al. (2022), the decomposition applies to all unsupervised methods except for Graph Factorization (Ahmed et al., 2013), which does not include a negative function $N$. Examples of popular decomposable loss functions are matrix reconstruction error (e.g., spectral embeddings) as well as softmax (e.g. node2vec (Grover and Leskovec, 2016) and LINE (Tang et al., 2015)). The decomposition also applies to supervised methods that regularize for graph structure ($\beta>0$ as defined in Chami et al. (2022)), such as Neural Graph Machines (Bui et al., 2018) and Planetoid (Yang et al., 2016).
@@
Given that graphs are sparse, performing gradient descent on $N$ is costly as the gradient update repels all dissimilar pairs, resulting in $\mathcal{O}\left(n^{2}\right)$ vector additions per epoch. In this paper, we build on the argument that the costly node-wise operation can be replaced with a more efficient dimension-wise operation, as summarized in Figure 1.
[…]
[Conceptual tree]A tree diagram summarizing our theoretical contribution. The root node is labeled ”negative function” and branches into ”squared loss” and ”skip-gram loss”. An example of squared loss is dimension-covariance regularization. With skip-gram loss, one descendant is SGNS while the other is our dimension-mean regularization. Each method is labeled with its computational complexity.
@@
Figure 1. Nearly all unsupervised and many supervised graph embedding loss functions define a “negative” function that repels embeddings of dissimilar nodes. We show that instead of repelling pairs of nodes (orange), which is costly, the negative function in the popular skip-gram (SG) loss can be approximated with a dimension-mean regularizer. The regularizer is efficient given that $d\ll n$. Our result establishing a connection between the skip-gram loss and dimension-mean regularization is novel. Runtime complexities are expressed in terms of the number of vector additions per epoch with details in Section 3.2.1.
@@
Below, we will map two dominant negative functions found in graph embedding algorithms to dimension regularizations. The two loss functions are spectral loss functions and skip-gram loss functions. The mapping of spectral loss functions to dimension covariance regularization in Section 3.1 parallels recent approaches to non-contrastive learning (Bardes et al., 2022; Garrido et al., 2022). Our novel contribution, detailed in Section 3.2, is a mapping from the skip-gram loss to a regularizer that induces origin-centered dimensions. Proof for all propositions below are included in Appendix A.